OpenStudy (anonymous):

The third term of a G.P is 27 and sixth term is 8 ; find the sum to infinite terms. . . ANSWER = a=243/4 , r=2/3 , S=729/4

OpenStudy (anonymous):

any solution for this one?

OpenStudy (anonymous):

let the terms are t1,t2,t3,t4,t5........................... for G.P series, a is the first term and r is the common ratio=t2/t1. here, n th term tn=ar^(n-1). so from the given data, t3=ar^(3-1)=27 and t6=ar^(6-1)=8 so t6/t3=ar^5/ar^2 =>8/27=r^(5-2) =>(2/3)^3=r^3 => r=2/3

OpenStudy (anonymous):

thanks for the r

OpenStudy (anonymous):

again t3=ar^2 =>a=t3/r^2=27/(2/3)^2=27*9/4=243/4

OpenStudy (anonymous):

and how do i get a and S

OpenStudy (anonymous):

now the sum for infinite series with r<1 is, s=a/(1-r)=(243/4)/(1-2/3)=(243/4)/(1/3)=243*3/4=729/4